OpenStudy (anonymous):

how do i solve (x^3y^-2/z^-5)^-4

OpenStudy (anonymous):

I'm thinking maybe y^8/(x^12 z^20)

OpenStudy (ivancsc1996):

\[(\frac{ x ^{3}y ^{-2} }{ z ^{-5} })^{-4}=\frac{ x ^{-12}y ^{8} }{ z ^{20} }=\frac{ y ^{8} }{ x ^{12}z ^{20} }\]